x ounces of pure acid is added to 105 ounces of 20% acid solution. How much should be added so the new mixture is 50% acid?
Answer by ankor@dixie-net.com(22740)

x ounces of pure acid is added to 105 ounces of 20% acid solution.
How much should be added so the new mixture is 50% acid?
:
.20(105) + x = .50(x+105)
21 + x = .5x + 52.5
x - .5x = 52.5 - 21
.5x = 31.5
multiply both sides by 2
x = 63 oz of pure acid required